Two blocks with masses 4.00 kg and 8.00 kg, respectively, are connected by a string and slide down a 30.0° inclined plane, as shown below. The coefficient of kinetic friction between the 4.00-kg block and the plane is 0.20; that between the 8.00-kg block and the plane is 0.40 (a) Calculate the acceleration of each block. (b) Calculate the tension in the string μ2-04 8 kg 4 kg 0-300


1
Expert's answer
2021-03-02T18:03:54-0500


a)


"n_1=m_1g\cos{30}\\f_{k1}=\mu_{k1}m_1g\cos{30}\\m_1a_1=m_1g\sin{30}-\mu_{k1}m_1g\cos{30}\\a_1=9.8(\sin{30}-0.2\cos{30})=3.20\frac{m}{s^2}"

"n_2=m_2g\cos{30}\\f_{k2}=\mu_{k2}m_2g\cos{30}\\m_2a_2=m_2g\sin{30}-\mu_{k2}m_2g\cos{30}\\a_2=9.8(\sin{30}-0.4\cos{30})=1.51\frac{m}{s^2}"

b)


"m_1a_1=m_1g\sin{30}-\mu_{k1}m_1g\cos{30}-T\\m_2a_2=m_2g\sin{30}-\mu_{k2}m_2g\cos{30}+T"

"T=\frac{4\cdot8(3.2-1.51)}{4+8}=4.51\ N"